Oakwood Elementary Fall Carnival gets bigger every year! Last year, they had 70 volunteers working 14 booths. This year, they want to have 42 booths.

If they use the same ratio of volunteers to booths, how many volunteers does Oakwood Elementary need this year?



1
Expert's answer
2021-11-16T17:54:28-0500

"\frac{70}{14}=\frac{x}{42}"

"x=\frac{70}{14}42=210" volunteers.